OpenStudy (anonymous):

Find the first term and the common ratio in a geometric progression, given that \(T_4-T_2=240\) and \(T_6-T_4=6000.\)

OpenStudy (helder_edwin):

since it is a geometric progression then \[ \large T_n=r^{n-1}T_1\qquad\text{for }n\geq2 \].

OpenStudy (helder_edwin):

then \[ \large 240=T_4-T_2=r^3T_1-rT_1=(r^3-r)T_1 \] and \[ \large 6000=T_6-T_4=r^5T_1-r^3T_1=(r^5-r^3)T_1 \]

OpenStudy (helder_edwin):

then \[ \large T_1=\frac{240}{r^3-r}\qquad\text{and}\qquad T_1=\frac{6000}{r^5-r^3} \]

OpenStudy (helder_edwin):

can u finish?

OpenStudy (anonymous):

Isn't \(T_n=ar^{n-1}?\)